OpenStudy (anonymous):

sin(arcsin(-11π/13) I'd really just like an explanation of this one.

OpenStudy (anonymous):

I don't know why I'm forgetting how to do this, but it's been awhile and my final is coming up and I want a better understanding of it.

OpenStudy (anonymous):

do you know the init Circle?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

when you at look the unit circle what quard would Sin be in?

OpenStudy (anonymous):

I see that it goes -11π, so I believe it's going clockwise around the unit circle and I think it gets to Q3.

OpenStudy (anonymous):

yes correct, so in the problem what does that tell you?

OpenStudy (anonymous):

I don't know...hahaha, rephrase that maybe?

OpenStudy (anonymous):

from looking at this problem and taking the information you already have what would you say the answer is? sin(arcsin(-11π/13)

OpenStudy (anonymous):

do you understand arcsin and what that means??

OpenStudy (anonymous):

crap, I wrote it down wrong. It was arcsin(sin(-11π/13))

OpenStudy (anonymous):

I have a pretty good understanding of how to use it. I'm rusty, though.

OpenStudy (anonymous):

okay, let us take this and break it down on what you know, you start I will fill in if needed.

OpenStudy (anonymous):

I believe sin(ø) = arcsin (-11π/13) in this situation.

OpenStudy (anonymous):

you are right, but how did you do that?

OpenStudy (anonymous):

I just kind of have that memorized. I guess I never really grasped that...

OpenStudy (anonymous):

okay, now I have a stupid what is the question asking for??? I am a little lost with the question here does it ask to solve or what?

OpenStudy (anonymous):

Find the exact value of the given expression. Simplify your answer (rationalize the denom)

OpenStudy (anonymous):

I also have that the answer is -2π/13

OpenStudy (anonymous):

I don't really understand where the 2π comes from.

OpenStudy (anonymous):

because recall the unit circle has a total of 360 degrees right?

OpenStudy (anonymous):

Yep and 2π is 360º

OpenStudy (anonymous):

2/pie is a point on the Unit the circle in which the angle is 360.. you beat to that one good job

OpenStudy (anonymous):

So, my picture shows the line -11π/13 and the distance between that line and the 180º line is 2π/13.

OpenStudy (anonymous):

Having a hard time understanding why it's 2π and not just π.

OpenStudy (anonymous):

Probably just overthinking it.

OpenStudy (anonymous):

what I am questioning here the negative here in this case?

OpenStudy (anonymous):

because you start with a negative you will most likely end that way..

OpenStudy (anonymous):

2pie is at the 360 degrees which has points 0,0 so how can that be a negative result help here?

OpenStudy (anonymous):

One sec, gonna post a photo of what is giving me a headache.

OpenStudy (anonymous):

I think we are both getting the same headache on this. This problem is going down..

OpenStudy (anonymous):

Hahaha, I'm tellin' you. I haven't had any problems with any other questions, but this one is just destroying my brain.

OpenStudy (anonymous):

yeah, the same thing only that my brain is dead because I am stuying Vectors myself.

OpenStudy (anonymous):

what I am missing here????

OpenStudy (lgbasallote):

is the question still \[\sin (arcsin (-11 \pi/3)\]

OpenStudy (anonymous):

arcsin(sin(-11π/3))

OpenStudy (anonymous):

yes sir, please help me I am getting a major headache here

OpenStudy (anonymous):

Messed up the on the original post

OpenStudy (lgbasallote):

here's a hint.. arcsin( sin x) = x sin (arcsin x) = x cos( arc cos x) = x arc cos (cos x) = x tan (arctan x) = x arctan (tan x) = x does that help?

OpenStudy (anonymous):

There is the picture that I was trying to describe.

OpenStudy (anonymous):

I wish that helped, but I still can't figure out why it is -2π/13

OpenStudy (lgbasallote):

is that included in the question? or is that your solytion?

OpenStudy (anonymous):

Solution. It's an answer key and I'm trying to decipher it.

OpenStudy (anonymous):

I know it is -2π/13 because of the range, but I don't know where the 2π came from.

OpenStudy (lgbasallote):

use the hint i gave you...here's a demo \[\sin (\sin^{-1} ( \frac{\pi}{2} ) ) = \frac{\pi}{2}\] \[\sin^{-1} (\sin ( \frac{2 \pi}{5} ) ) = \frac{2 \pi}{5}\] getting it now?

OpenStudy (lgbasallote):

\[\sin^{-1} (\sin ( \frac{-11\pi}{3} ) ) = -\frac{11\pi}{3}\] then simplify that you get \[-\frac{2\pi}{3}\] make sense?

OpenStudy (anonymous):

lgbasallote what did I miss here? I mean I thought for sure I was going the right way what did I do in my steps that went wrong here? Help me to understand..

OpenStudy (anonymous):

How'd you simplify to that?

OpenStudy (lgbasallote):

i think i wrote the wrong thingies lol...

OpenStudy (lgbasallote):

but that's the concept

OpenStudy (anonymous):

Yeah, I knew what you were saying with the others. It's this deal here. I don't understand how it is simplifying from 11π to 2π.

OpenStudy (anonymous):

It's like it just magically disappears.

OpenStudy (lgbasallote):

reference angles

OpenStudy (anonymous):

Heh, I feel dumb for not knowing still.

OpenStudy (anonymous):

Mrcarey I am so sorry I took the long way here.

OpenStudy (anonymous):

Haha, no problem at all. I've been at it for so long, it didn't even feel that long.

OpenStudy (anonymous):

ha ha yeah better next time right? HA HA..

OpenStudy (anonymous):

Exactly! Thanks for trying at least. Good luck with the vector calc. I'll be THERE soon enough.

OpenStudy (anonymous):

I am Physics Engineering major myelf

OpenStudy (anonymous):

Mech E. for me. Math math math math and more math!

OpenStudy (anonymous):

you got this now

OpenStudy (anonymous):

Yeah, I feel better about it. Got one more day of review in class, so I'll probably bring it up with him anyway. Thanks again!

OpenStudy (anonymous):

do you another problem you need help on? Just asking here

OpenStudy (anonymous):

Nope, this one seemed to be the only one I kept coming back to. This stupid thing made everything else seem easy!

OpenStudy (anonymous):

yeah, I only hope for me it is just one problem that will allow everything to be easy after that. Fat chance ha ha

OpenStudy (anonymous):

Hahahaha, no way. I've peeked at that stuff and needed to take a nap after.

OpenStudy (anonymous):

did that one already.. I took a 5hrs nap

OpenStudy (anonymous):

I am doing Calc1 andPhysics 1 with Calc

OpenStudy (anonymous):

I'll be doing that in the Fall. Well, no Physics yet, couldn't get in the class. General Chem 1 and Calc 1. Gonna be exciting.

OpenStudy (anonymous):

what school?

OpenStudy (anonymous):

Doing community for now, so I can transfer to University of Washington Seattle.

OpenStudy (anonymous):

You?

OpenStudy (anonymous):

cool I live Iin Portland Ore myself..

OpenStudy (anonymous):

I will start Sept 24th at Jr College than transfer to Tarleton ST university in TX in Spring 2013

OpenStudy (anonymous):

Awesome, TX should be pretty fun. Definitely a change of scenery.

OpenStudy (anonymous):

I am studying math and Physics by watching videos and taking notes on my own

OpenStudy (anonymous):

Khan Academy?

OpenStudy (anonymous):

yeah StephenVille is very small town though

OpenStudy (anonymous):

who is Khan A?

OpenStudy (anonymous):

www.khanacademy.com MIT grad who made his own site that teaches every kind of math and most sciences.

OpenStudy (anonymous):

Free site for learning. I've used it for awhile now. Especially when I get bad professors.

OpenStudy (anonymous):

funny you just posted MIT, Because I am using thier open courses on line for the video in math and Physics

OpenStudy (anonymous):

Haha, that works, too! You'd probably like his teaching style. He is very helpful and has tons of videos. I always recommend the site.

OpenStudy (anonymous):

thanks will check it out, hey you have a great night and if you need me to help me message me

OpenStudy (anonymous):

you

OpenStudy (anonymous):

Sounds like a plan! Have a good night!

OpenStudy (anonymous):

you too.